# Conditional Generative Modeling for Enhanced Credit Risk Management in Supply Chain Finance

### v1

@@ 6 Conclusion
This paper introduced a novel SCF risk management framework that leverages generative model to provide a full distributional understanding of sales, developing a unified analytical framework that simultaneously addresses credit risk assessment and loan size determination—objectives traditionally treated separately. By integrating QRGMM—a quantile-regression-based generative modeling method—with DeepFM to capture complex feature interactions, we have developed a flexible approach capable of estimating entire conditional distributions in complex SCF settings. This allows practitioners to move beyond conventional point estimations and simple risk classifications, offering a data-driven, distribution-oriented perspective on SCF risk.
@@
Building on the linear quantile regression framework of Hong et al. (2023), we establish rigorous asymptotic guarantees for three key risk measures, which serve as universal building blocks for deriving a wide range of risk metrics. Numerical experiments on synthetic data confirm that QRGMM accurately and stably approximates the underlying distribution, surpasses CWGAN in unconditional and conditional distribution matching, and recovers risk measures that align closely with known truths across the range of loan level space, greatly facilitating loan-sizing decision making in an intuitive way. On real-world datasets, QRGMM demonstrates practical applicability, accurately estimating quantiles and generating distributions that mirror real sales patterns.
@@
In summary, this work represents an initial exploration of the use of generative AI for operations management, advancing existing SCF risk management methodologies by employing generative models that provide richer distributional insights. However, much work remains to be done in this area. Future research should extend the framework to handle time-series data with seasonal effects and tackle challenges related to high-dimensional datasets. Additionally, while this study focuses on a single product for an SME, future efforts could explore loan portfolio problems by accounting for intricate interdependencies among products, including complementarity and substitution effects. And applications in other domains such as financial risk management and supply chain optimization could also be investigated. By continuing to refine these generative distributional approaches, the ultimate goal is to support more informed, resilient, and data-driven decision-making in various fields of operations management.

### v3

@@ 7 Conclusion
This paper introduced a novel SCF risk management framework that leverages a generative model to provide a full distributional understanding of sales, developing a unified analytical framework that simultaneously addresses credit risk assessment and loan size determination, objectives traditionally treated separately. By integrating QRGMM, a quantile-regression-based generative modeling method, with DeepFM to capture complex feature interactions, we have developed a flexible approach capable of estimating entire conditional distributions in complex SCF settings. This allows practitioners to move beyond conventional point estimation and simple risk classification methods, offering a data-driven, distribution-oriented perspective on SCF risk.
@@
Building on the linear quantile regression framework of hong2023learning, we establish rigorous asymptotic guarantees for the key risk measures, which serve as universal building blocks for deriving a wide range of risk measures. Numerical experiments on synthetic data confirm that QRGMM accurately and stably approximates the underlying distribution, surpasses CWGAN in unconditional and conditional distribution matching, and recovers risk measures that align closely with true benchmarks across the range of loan level space, greatly facilitating loan-sizing decision making in an intuitive way. On real-world datasets, QRGMM demonstrates practical applicability, accurately estimating quantiles, generating distributions that mirror real sales patterns, and accurately estimating risk measures.
@@
In summary, this work represents an initial exploration of the use of generative models for operations management, advancing existing SCF risk management methodologies by employing generative models that provide richer distributional insights. However, much work remains to be done in this area. Future research should extend the framework to handle time-series data with seasonal effects and tackle challenges related to high-dimensional datasets. Additionally, while this study mainly focuses on a single product for a seller, future efforts could explore loan portfolio problems by accounting for intricate interdependencies among products, including complementary and substitutable relationships. Moreover, applications in other domains, such as financial risk management and supply chain optimization, could also be investigated. By continuing to refine these generative distributional approaches, the ultimate goal is to support more informed, resilient, and data-driven decision-making in various fields of operations management.
Appendix A Proof of Proposition 1
Proof.
Recall that $\tau_{j}=\frac{j}{m}$ with $\tau_{j+1}-\tau_{j}=\frac{1}{m}$. For any $j\in\{1,\ldots,m-1\}$ such that $\tau_{j},\tau_{j+1}]\cap[\tau_{\mathsf{L}}+\frac{1}{m},\tau_{\mathsf{U}}-\frac{1}{m}]\neq\emptyset$, let $\tau\in[\tau_{j},\tau_{j+1}]\cap[\tau_{\mathsf{L}}+\frac{1}{m},\tau_{\mathsf{U}}-\frac{1}{m}]$. For any given $\mathbf{x}$, define
$\tilde{\alpha}=\frac{\hat{Q}(\tau_{j+1}\mid\mathbf{x})-\hat{Q}(\tau\mid\mathbf{x})}{\hat{Q}(\tau_{j+1}\mid\mathbf{x})-\hat{Q}(\tau_{j}\mid\mathbf{x})},\qquad 1-\tilde{\alpha}=\frac{\hat{Q}(\tau\mid\mathbf{x})-\hat{Q}(\tau_{j}\mid\mathbf{x})}{\hat{Q}(\tau_{j+1}\mid\mathbf{x})-\hat{Q}(\tau_{j}\mid\mathbf{x})},$
and because $F_{\hat{Y}}\left(\hat{Q}(\tau\mid\mathbf{x})\mid\mathbf{x}\right)$ is linearly interpolated between $\tau_{j}=F_{\hat{Y}}\left(\hat{Q}(\tau_{j}\mid\mathbf{x})\mid\mathbf{x}\right)$ and $\tau_{j+1}=F_{\hat{Y}}\left(\hat{Q}(\tau_{j+1}\mid\mathbf{x})\mid\mathbf{x}\right)$, we have $F_{\hat{Y}}\!\left(\hat{Q}(\tau\mid\mathbf{x})\mid\mathbf{x}\right)=\tilde{\alpha}\,\tau_{j}+(1-\tilde{\alpha})\,\tau_{j+1}.$ Then
$F_{\hat{Y}}\!\left(\hat{Q}(\tau\mid\mathbf{x})\mid\mathbf{x}\right)-F_{Y}\!\left(\hat{Q}(\tau\mid\mathbf{x})\mid\mathbf{x}\right)$
$=\tilde{\alpha}\bigl[\tau_{j}-F_{Y}(\hat{Q}(\tau_{j}\mid\mathbf{x})\mid\mathbf{x})\bigr]+(1-\tilde{\alpha})\bigl[\tau_{j+1}-F_{Y}(\hat{Q}(\tau_{j+1}\mid\mathbf{x})\mid\mathbf{x})\bigr]$
$\quad+\tilde{\alpha}\!\left\{F_{Y}(\hat{Q}(\tau_{j}\mid\mathbf{x})\mid\mathbf{x})-F_{Y}(\hat{Q}(\tau\mid\mathbf{x})\mid\mathbf{x})\right\}+(1-\tilde{\alpha})\!\left\{F_{Y}(\hat{Q}(\tau_{j+1}\mid\mathbf{x})\mid\mathbf{x})-F_{Y}(\hat{Q}(\tau\mid\mathbf{x})\mid\mathbf{x})\right\}.$
Using the identity
$F_{Y}(\hat{Q}(\tau_{j+1}\mid\mathbf{x})\mid\mathbf{x})-F_{Y}(\hat{Q}(\tau_{j}\mid\mathbf{x})\mid\mathbf{x})=\bigl[F_{Y}(\hat{Q}(\tau_{j+1}\mid\mathbf{x})\mid\mathbf{x})-\tau_{j+1}\bigr]+\tfrac{1}{m}+\bigl[\tau_{j}-F_{Y}(\hat{Q}(\tau_{j}\mid\mathbf{x})\mid\mathbf{x})\bigr],$
and the triangle inequality, we obtain the bound
$\bigl|F_{\hat{Y}}(\hat{Q}(\tau\mid\mathbf{x})\mid\mathbf{x})-F_{Y}(\hat{Q}(\tau\mid\mathbf{x})\mid\mathbf{x})\bigr|$ (5)
  $\leq 3\max\!\Bigl\{\bigl|\tau_{j}-F_{Y}(\hat{Q}(\tau_{j}\mid\mathbf{x})\mid\mathbf{x})\bigr|,\;\bigl|\tau_{j+1}-F_{Y}(\hat{Q}(\tau_{j+1}\mid\mathbf{x})\mid\mathbf{x})\bigr|\Bigr\}+\frac{1}{m}.$
By Assumption (2.a), for each grid point $\tau_{j}$ there exists $\eta_{j}$ between $Q(\tau_{j}\mid\mathbf{x})$ and $\hat{Q}(\tau_{j}\mid\mathbf{x})$ such that
$\bigl|\tau_{j}-F_{Y}(\hat{Q}(\tau_{j}\mid\mathbf{x})\mid\mathbf{x})\bigr|=\bigl|F_{Y}(Q(\tau_{j}\mid\mathbf{x})\mid\mathbf{x})-F_{Y}(\hat{Q}(\tau_{j}\mid\mathbf{x})\mid\mathbf{x})\bigr|=f_{Y}(\eta_{j}\mid\mathbf{x})\,\bigl|Q(\tau_{j}\mid\mathbf{x})-\hat{Q}(\tau_{j}\mid\mathbf{x})\bigr|.$
Let
$\underline{f}:=\inf_{\tau\in[\tau_{\ell},\tau_{u}]}\inf_{\mathbf{x}\in\mathcal{X}}f_{Y}(F_{Y}^{-1}(\tau\mid\mathbf{x})\mid\mathbf{x})>0,\qquad\overline{f}:=\sup_{\tau\in[\tau_{\ell},\tau_{u}]}\sup_{\mathbf{x}\in\mathcal{X}}f_{Y}(F_{Y}^{-1}(\tau\mid\mathbf{x})\mid\mathbf{x})<\infty,$
as guaranteed by Assumption (2.b). Then, for all $j$ with $\tau_{j}\in[\tau_{\ell},\tau_{u}]$,
$\bigl|\tau_{j}-F_{Y}(\hat{Q}(\tau_{j}\mid\mathbf{x})\mid\mathbf{x})\bigr|\leq\overline{f}\;\bigl|Q(\tau_{j}\mid\mathbf{x})-\hat{Q}(\tau_{j}\mid\mathbf{x})\bigr|.$
Hence,
$\max_{\tau_{j}\in[\tau_{\ell},\tau_{u}]}\bigl|\tau_{j}-F_{Y}(\hat{Q}(\tau_{j}\mid\mathbf{x})\mid\mathbf{x})\bigr|\;\leq\;\overline{f}\;\max_{\tau_{j}\in[\tau_{\ell},\tau_{u}]}\bigl|Q(\tau_{j}\mid\mathbf{x})-\hat{Q}(\tau_{j}\mid\mathbf{x})\bigr|.$ (6)
By Lemma 2,
$\max_{\tau_{j}\in[\tau_{\ell},\tau_{u}]}\bigl|Q(\tau_{j}\mid\mathbf{x})-\hat{Q}(\tau_{j}\mid\mathbf{x})\bigr|\;=\;O_{\mathbb{P}}\!\left(\frac{1}{\sqrt{n}}\right).$
Combining this with (5)–(6) and taking the supremum over $\tau\in[\tau_{\ell}+1/m,\tau_{u}-1/m]$ yields
$\sup_{\tau\in[\tau_{\ell}+1/m,\tau_{u}-1/m]}\bigl|F_{\hat{Y}}(\hat{Q}(\tau\mid\mathbf{x})\mid\mathbf{x})-F_{Y}(\hat{Q}(\tau\mid\mathbf{x})\mid\mathbf{x})\bigr|\;=\;O_{\mathbb{P}}\!\left(n^{-1/2}\right)+O\!\left(m^{-1}\right).$
Finally, since $y=\hat{Q}(\tau\mid\mathbf{x})$ ranges over the corresponding support $\mathcal{Y}_{m,n}(\mathbf{x})$, we conclude
$\sup_{y\in\mathcal{Y}_{m,n}(\mathbf{x})}\bigl|F_{\hat{Y}}(y\mid\mathbf{x})-F_{Y}(y\mid\mathbf{x})\bigr|\;=\;O_{\mathbb{P}}\!\left(n^{-1/2}\right)+O\!\left(m^{-1}\right),$
Appendix B Sensitivity of $m$
We examine the sensitivity of QRGMM with respect to the number of quantile grid points $m$ using the synthetic FM location–scale shift model from Section 6.1. For each replication, the model was trained under different values of $m$ (logarithmically spaced), with training data sizes n=8,000n=8{,}000 and n=2,800n=2{,}800, corresponding to the training data sizes used in Sections 6.1 and 6.3, respectively. Evaluation was conducted on 10,00010{,}000 observations of $\hat{Y}(\mathbf{x})$ generated by QRGMM at a fixed covariate $\mathbf{x}$, where the Kolmogorov–Smirnov statistic was computed against the true distribution. Each experiment was repeated for 50 replications, and the results were averaged over these replications, denoted by $\bar{D}_{m}$. Figure 23 displays $\bar{D}_{m}$ as a function of $m$ for QRGMM.
The convergence rate analysis in Section 4.3 shows that the error of the linear QRGMM follows the rate $O_{\mathbb{P}}(n^{-1/2})+O(m^{-1})$, suggesting $m$ should be chosen on the order of $\sqrt{n}$. Our experiments indicate that, after incorporating the deep neural network and factorization machine components in QRGMM, the performance is not highly sensitive to the choice of $m$. Hence, we adopt $m=\sqrt{n}$ as a robust and convenient choice throughout the paper.
(a) $n=8000$
(b) $n=2800$
Figure 23: Sensitivity of QRGMM to the number of quantile grid points $m$ under different training sample sizes.
Appendix C Effects of Covariates Changes on Conditional Distributions and Loan Decisions
We use the conditional test evaluations in Section 6.1 to illustrate how covariates influence conditional distributions and subsequent loan sizing decisions. Specifically, we choose two covariate vectors randomly drawn from a uniform distribution over the covariate space to conduct the conditional test. In Section 6.1, the conditional test is conducted at
[…]
while here we consider
[…]
Although both $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ are randomly drawn, they represent substantially different feature combinations, enabling us to examine how such shifts affect the conditional distribution of $Y(\mathbf{x})$ and the resulting loan sizing decision.
For this new covariate $\mathbf{x}_{2}$, we repeat the conditional test evaluations described in Section 6.1. Specifically, we report the sample means and standard deviations of generated conditional datasets (Table 4), together with Wasserstein distance (WD) and Kolmogorov–Smirnov statistics (KS), averaged over 100 replications (Figures 24 and 25). Visual comparisons between generated and true conditional distributions (Figure 26), as well as the estimation of risk measures $\{r_{1}(l),r_{2}(l),\text{LGD}(l),r_{3}(l)\}$ (Figure 27), are also presented below.
Table 4: Mean and standard deviation (SD) of generated samples.
Model Unconditional Test Conditional $Y(\mathbf{x}_{1})$ Conditional $Y(\mathbf{x}_{2})$
Mean SD Mean SD Mean SD
Truth $4355.4783$ $1430.8676$ $3948.9500$ $150.6615$ $3832.1420$ $148.9017$
QRGMM $4350.6580$ $1412.5260$ $3948.7998$ $150.1433$ $3814.0224$ $145.0772$
QRGMM $4355.9730$ $1421.0893$ $4114.5597$ $221.1148$ $3867.7647$ $165.1972$
CWGAN $4320.6589$ $1394.2039$ $3915.1583$ $190.8393$ $3761.0674$ $141.0797$
(a) $\mathbf{x}_{1}$
(b) $\mathbf{x}_{2}$
Figure 24: WD for conditional test under different covariates.
(a) $\mathbf{x}_{1}$
(b) $\mathbf{x}_{2}$
Figure 25: KS for conditional test under different covariates.
(a) $\mathbf{x}_{1}$
(b) $\mathbf{x}_{2}$
Figure 26: Histograms of generated vs. true data under different covariates.
(a) $\mathbf{x}_{1}$
(b) $\mathbf{x}_{2}$
Figure 27: Estimation of risk measures across loan levels $l$ under different covariates.
Summary of Results.
The results clearly show that the conditional distribution shifts when covariates change, leading to different loan sizing outcomes. For example, the mean and standard deviation of the conditional distribution under $\mathbf{x}_{2}$ differ from those observed under $\mathbf{x}_{1}$ (see Table 4), which in turn alters the estimated default probabilities and loss measures (see Figure 27). Nevertheless, QRGMM continues to deliver stable and accurate estimation: its generated observations align closely with the true conditional distribution across all metrics, and its risk measure curves consistently track the analytical benchmarks. In contrast, both QRGMM and CWGAN exhibit unstable and poor performance, indicating notable limitations in their effectiveness. These findings underscore both the sensitivity of loan decisions to covariates changes and the robustness of QRGMM across these variations.